A fullback with a mass of 100 kg and a velocity of 3.0 m/s due west collides head-on with a defensive back with a mass of 86 kg and a velocity of 4.5 m/s due east. (Take the positive direction to be to the west.) (a) What is the initial momentum of each player? momentum of full back kg · m/s momentum of defensive back kg · m/s (b) What is the total momentum of the system before the collision? kg · m/s (c) If they stick together and external forces can be ignored, what direction will they be traveling immediately after they collide? east west
Answer #1

Solution:

a) Initial momentum of fullback = m * v

P_fullback = 100 * 3 = 300 kg m/s

Initial momentum of defensive back

=> P_defensive back = 86 * (-4.5) = -387 kg m/s

b) Total momentum before collision

=> P_total = P_fullback + P_defensive back

=> P_total = 300 + (-387)

=> P_total = -87 kg m/s

c) As the total momentum is negative hence if the stick together they will move east.

east is the right answer.

thanks.